### v1

# One-Shot Graph Representation Learning Using Hyperdimensional Computing

Abstract
We present a novel, simple, fast, and efficient approach for semi-supervised learning on graphs. The proposed approach takes advantage of hyper-dimensional computing which encodes data samples using random projections into a high dimensional space (HD space for short). Specifically, we propose a Hyper-dimensional Graph Learning (HDGL) algorithm that leverages the injectivity property of the node representations of a family of graph neural networks. HDGL maps node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. Results of experiments with widely used benchmark data sets show that HDGL achieves predictive performance that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive training.
1 Introduction
Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations. Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction Li and Pi (2020); Kumar et al. (2020). Such problems are typically solved using graph embedding techniques Zhou et al. (2022); Cai et al. (2018); Xu (2021); Wang et al. (2022).
Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita (1997); Frasconi et al. (1998); Gori et al. (2005); Scarselli et al. (2009), which learn low-dimensional embeddings of nodes in a graph by aggregrating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning Kipf and Welling (2017); Veličković et al. (2018); Hamilton et al. (2017), because of their superior performance across a broad spectrum of applications Wu et al. (2020); Zhang et al. (2018); Zhou et al. (2022).
Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact Desislavov et al. (2023); Strubell et al. (2020); de Vries (2023); Thompson et al. (2021). Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
Hyperdimensional computing Kanerva (1988, 2009); Neubert et al. (2019); Thomas et al. (2021) offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector Kanerva (1988, 2009). All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al. (2019); Thomas et al. (2021). Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware Kleyko et al. (2023a, b). Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning Neubert et al. (2019); Kleyko et al. (2023b); Ge and Parhi (2020); Morris et al. (2022).
Against this background, we investigate Hyper-dimensional Graph Learner (HDGL), a one-shot approach to graph learning, that makes use of hyper-dimensional computing Kanerva (1988, 2009); Neubert et al. (2019); Thomas et al. (2021). Specifically, we outline a methodology for designing an HD mapping for node embedding that achieves competitive performance in semi-supervised node labeling tasks when compared to GNNs, which necessitate considerably more costly training. The key contributions of the paper are as follows:
  • We show that how to design a HD node embeddings that match the expressive power of GNN embeddings.
  • We propose HDGL, a graph learner that takes advantage of carefully designed hyperdimensional node embeddings to perform one-shot learning on graphs for semi-supervised node labeling tasks.
  • We present results of extensive experimental comparisons of HDGL with the state-of-the-art GNN methods on several standard benchmark data sets which show that HDGL achieves performance that is competitive with that of GNN, while avoiding the need for computationally expensive training.
The rest of the paper is organized as follows: Section 2 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN. Section 3 introduces HDGL, the proposed HD computing based one-shot algorithm for learning node embeddings. Section 4 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods. Section 5 concludes with a brief summary of the paper and an outline of some directions for further research.
2 Background and Preliminaries
2.1 Node Embedding Problem
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. Let $y_{v}\in\{1,\cdots,L\}$ be the the label associated with each $v\in\mathcal{V}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$. The goal is to learn a latent representation $\phi(v)$ of each node $v$ using the structure of $\mathcal{G}$ and $\mathbf{x}_{v}$ to accurately predict $y_{v}$.
2.2 Properties of HD Computing
Hyper-dimensional computing, originally introduced by Kanerva (1988) is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional static mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large Kanerva (1988, 2009). All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al. (2019). The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations Kleyko et al. (2023a, b). In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$. We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ Kanerva (1988). This is also known as a blessing of HD spaces.
High Memory Capacity of HD Representation. As noted by Neubert et al. (2019), in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits Ahmad and Hawkins (2015); Neubert et al. (2019); Thomas et al. (2021). Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms Neubert et al. (2019).
2.3 HD Computing Operations
Bundling, denoted by $\oplus$, is a bitwise majority operation represented by :$\oplus:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. Bundling two randomly sampled HD vectors results in a new vector which is similar to the input vectors. Bundling $\oplus$ is associative and commutative i.e $\mathbf{a}\oplus(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\oplus\mathbf{b})\oplus\mathbf{c}=(\mathbf{c}\oplus\mathbf{a})\oplus\mathbf{b}$. Note that in case of Bundling an even number of HD vectors, there is a possibility of ties since Bundling is a bitwise majority operation. To break ties, a tie-breaking policy, e.g., random choice of 0 or 1 on dimensions where ties occur. Bundling can be thought of as the HD space analog of the mean operation euclidean spaces. Bundling, because it commutes, can be used to represent sets as well as multisets of objects encoded using HD vectors.
Binding, denoted by $\otimes$, is a bitwise Exclusive OR operation represented by :$\otimes:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, is an operation which takes two input vectors and maps them to a new vector. An important property of Binding is that Binding two HD vectors results in a vector that is dissimilar to both. Formally, let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two sampled HD vectors and $\mathbf{c}=\mathbf{a}\otimes\mathbf{b}$. Due to the near-orthogonality of HD vectors, we can then say that $d_{H}[\mathbf{c},\mathbf{a}]\approx 0.5$ and $d_{H}[\mathbf{c},\mathbf{b}]\approx 0.5$. Binding is invertible, associative and commutative i.e $\mathbf{b}\otimes(\mathbf{a}\otimes\mathbf{b})=(\mathbf{b}\otimes\mathbf{a})\otimes\mathbf{b}=\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$. Moreover, Binding is distributive over bundling i.e $\mathbf{a}\otimes(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\otimes\mathbf{b})\oplus(\mathbf{a}\otimes\mathbf{c})$. Note that Binding also has a reflective property i.e the distance between two vectors doesn’t change when both vectors are bound to the same vector. Formally, $d_{H}[\mathbf{a},\mathbf{b}]=d_{H}[\mathbf{c}\otimes\mathbf{a},\mathbf{c}\otimes\mathbf{b}]$. Note that this also works even vectors are slightly different i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\mathbf{c}\otimes\mathbf{a}^{\prime},\mathbf{c}\otimes\mathbf{b}]$ if $\mathbf{a}\approx\mathbf{a}^{\prime}$.
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler (1998) in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
Note that operations and vectors mentioned above have counterparts in the bipolar space $\{-1,1\}^{n}$.
2.4 GNN-based Node Embdeddings
Recall that GNN use the node features, together with the adjacency information to learn the embedding $\mathbf{h}_{v}$ of each node $v\in\mathcal{V}$. Most GNN use a neighborhood aggregation scheme where each node recursively aggregates the features of its neighbors to compute its feature vector Xu et al. (2019); Wu et al. (2020); Zhang et al. (2018); Zhou et al. (2022). After $k$ iterations of such aggregation, the node’s feature vector yields its embedding. More precisely, the latent representation of node $v$ at the $k$th iteration is given by:
$\mathbf{h}^{(k)}_{v}=\text{Agg}(\{\mathbf{W}^{(k)}\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\cup\{v\}\})$
[…]
2.5 Representational Power of GNN
Several authors have investigated the representational power of GNN from different perspectives, including their ability to distinguish between different graph structures Xu et al. (2019), e.g., using the Weisfeiler-Lehman (WL) test Weisfeiler and Leman (1968); Huang and Villar (2021) (which is known to decide graph isomorphism in general, barring a few exceptions, e.g., regular graphs), their ability to learn a set of functions on graphs Scarselli et al. (2009), and the equivalence of the preceding two perspectives Chen et al. (2019). The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors Weisfeiler and Leman (1968); Babai and Kucera (1979). The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors Huang and Villar (2021). Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into different representations. As shown by Xu et al. (2019); Morris et al. (2019), GNN are at most as powerful as the WL test in distinguishing graph structures; and for a GNN with a sufficient number of layers will be as powerful as the WL test if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme
$\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$
where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations Xu et al. (2019).
3 Hyper-Dimensional Graph Learner
In essence, our design of HDGL aims to achieve a high level of expressive power similar to graph neural networks (GNN). We utilize HD computing operations on node representations, taking advantage of HD operations—Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) —and their properties. This ensures that HDGL, akin to GNN, not only aspires to build expressive power but also learns to map similar graphs to similar representations.
The proposed HD learner carries out 3 major steps:
  1. i
    Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$.
  2. ii
    Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$.
  3. iii
    Finally, use $\mathbf{z}_{v}$ to predict $y_{v}$
We proceed to describe each of these steps in detail.
Figure 1: An example mapping node features to a binary space using Random Hyperplanes from Efimov (2023). A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side. Utilizing multiple hyperplanes results in a binary sketch for each point.
3.1 Mapping Node Features to HD-space
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations Dirksen et al. (2022). The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$.
Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector. Nodes with similar features will have similar HD representations; and nodes with dissimilar features will have dissimilar HD representations.
3.2 Neighborhood Aggregation Scheme
Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows:
$\begin{split}\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots\\ \ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}\end{split}$
where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
3.2.1 Bundling for generalizing across similar nodes
We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets). Hence, we propose to use Bundling to realize $\varphi_{k}$. More precisely,
$\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$
where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective. However, as noted by Xu et al. (2019), GNNs that use the mean to aggregate features over node neighborhoods Kipf and Welling (2017); Hamilton et al. (2017) perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
It is worth noting that the Binding operation $\otimes$ would be inappropriate in this context because binding of two similar vectors results in dissimilar vectors. Thus, if Binding were used to realize $\varphi_{k}$, similar yet distinct local neighborhoods structures would yield different HD representations, adversely impacting HDGL’s ability to generalize across such graph structures. To complicate matters, Binding is also invertible i.e $\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$; and hence, it could discard information that is useful for downstream task e.g., node label prediction.
3.2.2 Binding for linking node with its neighborhood
Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
We propose to use Binding the operation $\otimes$ to realize the function $\phi(\cdot)$ because it offers a means to balance the discriminative power of the resulting model against its generalization ability. We illustrate the preceding claim with an example. Let $v\in\mathcal{V}$ and $w\in\mathcal{V}$ be two nodes in a graph. Suppose $\phi(\cdot)$ is realized using the Binding Operation $\otimes$ and $\psi_{k}(\cdot)$ is realized using the Bundling Operation. Considering only 1-hop and 2-hop neighbors of $v$, the latent representation of $v$ and $w$ will be
$\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$ (1)
Suppose we have
$\begin{split}\mathbf{r}_{w}\approx\mathbf{r}_{v};\quad\;\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\approx\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\\ \end{split}$
and ${\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}}$ is dissimilar to ${\mathop{\oplus}_{j\in\mathcal{N}^{2}(w)}\mathbf{r}_{j}}$. Then $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since binding is reflecting i.e
$d_{H}\Big{[}\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)},\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(w)}\mathbf{r}_{j}\Big{)}\Big{]}\approx d_{H}\Big{[}\mathbf{z}_{v},\mathbf{z}_{w}\Big{]}$
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired.
On the other hand, if the terms representing the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
Conversely, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$. Bundling of the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$, a vector that has already been bundled with $k$-hop neighbors, will result in a vector that predominantly represents information from a node’s neighborhood, while leaving out essential information about the itself. This is due to the fact that Bundling is a bitwise majority operation; and hence, if bundling is used to realize $\phi(\cdot)$, the information from a node’s neighborhood suppresses the information about the node itself.
Dataset Nodes Edges Features Classes Label rate Edge density
CORA 2708 5429 1433 7 0.0563 0.0004
CiteSeer 3327 4732 3703 6 0.0569 0.0004
PubMed 19717 44338 500 3 0.0030 0.0001
Coauthor CS 18333 81894 6805 15 0.0164 0.0001
Coauthor Physics 34493 247962 8415 5 0.0029 0.0001
Table 1: Dataset Summary.
3.2.3 Improving injectivity via Rotation
Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as
$\begin{split}\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm. To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information. To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
where $\Pi(\cdot)$ and $\Pi\;\Pi(\cdot)$ denote single and double rotation operations respectively. The preceding equation ensures that $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since rotation of a HD vector yields a HD vector that is dissimilar/orthogonal to the original. That is,
$\begin{split}d_{H}\Big{[}\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)},\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\Big{]}\approx 0.5\\ d_{H}\Big{[}\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\big{)},\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\Big{]}\approx 0.5\\ \end{split}$
Hence, considering only 1-hop and 2-hop neighbors, the proposed aggregation scheme to be used by HDGL for computing the HD representation of a node $v\in\mathcal{V}$ in the graph is given by
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$ (2)
3.3 Node Label Prediction in HDGL
With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects Ge and Parhi (2020). Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD are constructed for each class as follows:
$\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$
resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows:
$\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}].$
4 Experiments
We proceed to report results of extensive experiments comparing HDGl with against a set of state-of-the-art node labeling methods, including several GNN models, in terms of both accuracy and runtime.
4.1 Data sets
For our experiments, we use the three of the most commoly used node labeling benchmarks: CORA McCallum et al. (2000), CiteSeer Giles et al. (1998), PubMed Namata et al. (2012) data sets and two data sets from the the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. (2018). We used the Train/Validation/Test splits described inYang et al. (2016); Kipf and Welling (2017); Veličković et al. (2018) for CORA, CiteSeer and PubMed data sets and that described in Shchur et al. (2018) for Coauthor CS and Coauthor Physics data sets. The data sets are summarized in Table 1.
4.2 Models
We compare the performance of HDGL with that of several strong baselinesL Multilayer Perceptron (MLP), DeepWalk Perozzi et al. (2014), Label Propagation. We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) Kipf and Welling (2017), Graph Attention Network (GAT) Veličković et al. (2018), and lastly, an fast GNN, namely, Simplified Graph Convolutional Network (SGC) Wu et al. (2019) which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
4.3 HDGL Model Specifications
To ensure that HDGL are directly comparable with most GNN architectures which typically use 2 layers, and hence aggregate information over 2-hop neighbors, we limit HDGL to aggregate information over at most 2-hops i.e HD node representations are calculated as per Eq.(2).
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL.
In the case of CORA and CiteSeer datasets, we map the node features to $50,000$-dimensional HD vectors via random projections. For the PubMed, Coauthor CS, and Coauthor Physics datasets, we map the node features $20,000$-dimensional HD vectors in a similar fashion. While these choices may seem arbitrary, we note that a choice between $20000$ and $50000$ is typically large enough for the desired properties of HD-computing to hold.
4.4 Training and Evaluation
Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters. Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance. In contrast, one-shot learning methods such as HDGL have no need for validation data. Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question. Thus, HDGL uses training and validation sets together to build the model, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning. In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. For tuning the hyperparameters, we used grid search over a set of 12 configurations. In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay. In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored. Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4]. In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6]. GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3. The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three models. As MLP, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. Instead, we choose to utilize publicly released implementations that already incorporate optimal hyperparameters.
All of the models were evaluated on the same test data. In each case, we report measured performance on the test data, averaged over 10 random choices of the relevant parameters: random hyperplanes used to construct HD representations used by HDGL, weight initializations of the deep learning models.
Method Cora Citeseer Pubmed CompSci Physics
MLP 55.1 $\pm$ 0.6 46.5 $\pm$ 0.4 71.4 $\pm$ 0.5 87.9 $\pm$ 0.6 88.2 $\pm$ 1.0
DeepWalk 67.2 $\pm$ 0.8 43.2 $\pm$ 0.6 65.3 $\pm$ 0.5 78.1 $\pm$ 0.9 87.6 $\pm$ 0.9
Label Prop 68.2 $\pm$ 1.1 45.3 $\pm$ 0.8 63.1 $\pm$ 0.5 74.1 $\pm$ 3.9 86.1 $\pm$ 2.1
GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 91.1 $\pm$ 0.8 92.8 $\pm$ 0.5
GAT 82.8 $\pm$ 0.5 71.0 $\pm$ 0.6 79.1 $\pm$ 0.3 90.5 $\pm$ 0.6 92.5 $\pm$ 0.9
SGC 79.8 $\pm$ 0.0 68.5 $\pm$ 0.1 75.4 $\pm$ 0.0 89.1 $\pm$ 0.1 92.1 $\pm$ 0.1
HDGL 79.5 $\pm$ 0.7 70.7 $\pm$ 0.6 77.8 $\pm$ 1.1 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9
Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of HDGL Hyperplanes Across All Datasets
4.5 Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework Paszke et al. (2019), augmented with the Deep Graph Library (DGL) Wang et al. (2019) for importing datasets and implementing deep learning models. None of our experiments used parallel execution or GPUs.
Figure 2: Comparison of Learning Times (in seconds) Across Datasets for HD-GL with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
4.6 Results
4.6.1 Predictive Performance
The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL, an one-shot learning method without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across all 5 benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (MLP, DeepWalk, LabelProp).
4.6.2 Run-times
Fig 2 compares run-time of HDGL (which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyper-parameter tuning) with the run-time of HDGL. We observe that barring one exception, PubMed data where the run-times of SGC and HDGL are close, the run-time of HDGL is substantially less than that of all the state-of-the-art GNN.
The speedup obtained by HDGL over the GNN methods is particularly noticeable for larger graph datasets, such as Coauthor CS and Physics. Figure 3, shows how the methods compare with respect to both predictive accuracy achieved and the run-time used to achieve it. It is apparent that HDGL achieves a substantial speedup by factors of 3.5x to 4 with only minimal (1-2%) drop in performance compared to GNN models.
In practice, deep learning models are often optimized using a far more extensive search over the space of hyperparameters, with corresponding increases in run-time, relative to HDGL, with the use of GPUs further widening the gap between GNN and HDGL in terms of their energy footprints.
Figure 3: Tradeoff of Predictive performance against Run-time for HDGL and GNN on Coauthor CS and Coauthor Physics Datasets. Note that GNN models (GCN, GAT, SGC) performance shown is achieved with hyperparameter tuning whereas HDGL offers one-shot learning without the need for hyperparameter tuning
5 Conclusion
We have presented a novel simple, one-shot, and hence fast, and efficient approach to semi-supervised algorithm for graph representation learning. The proposed approach constructs low precision, noise-tolerant, similarity preserving randomized hyper-dimensional representations of nodes in graphs. We take advantage of the close connection between the Weisfeiler-Lehman test for graph isomorphism and a family of graph neural networks to design a hyper-dimensional graph learning algorithm. The proposed algorithm uses hyper-dimensional operators such as bundling and binding to aggregate information from the local neighborhood of each node so as to represent nodes using information that is an approximation of that used by the Weisfeiler-Lehman test.
Results of experiments with widely used benchmark data sets show that the proposed algorithm achieves performance on semi-supervised node labeling tasks that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive optimization of model parameters and hyper-parameters. Consequently, the proposed method offers an especially attractive alternative with energy footprint that is substantially smaller than that of deep learning based modern graph embedding methods.
Because the model supports one-shot learning of node embeddings, it lends itself to use in settings that call for continual learning algorithms De Lange et al. (2021); Wickramasinghe et al. (2023) wherein data samples, or classes, are presented on an ongoing basis, allowing the model to be continually updated, without the need for retraining on old data. Additional directions for further research include variants of the proposed algorithm that can learn graph representations that invariant or equivariant with respect to a specified set of operations on graphs Satorras et al. (2021); Keriven and Peyré (2019).