### v2

# Hyperdimensional Computing for Node Classification and Link Prediction

Abstract
We introduce a novel method for transductive learning on graphs using hyperdimensional representations. The proposed approach encodes data samples using random projections into a very high-dimensional space (hyperdimensional or HD space for short). It obviates the need for expensive iterative training of the sort required by deep learning methods. Specifically, we propose a Hyperdimensional Graph Learning (HDGL) algorithm. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost. Furthermore, HDGL is well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes. Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods, although it falls short of state-of-the-art Graph Neural Network (GNN) methods that rely on computationally expensive iterative training. We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.
1 Introduction
Node Classification and Link Prediction in Networks. Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations. Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction [30, 29]. Such problems are typically solved using graph embedding techniques [63, 6, 60, 55].
1.1 Related Work
Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita [46], Frasconi et al. [13], Gori et al. [17], Scarselli et al. [44], which learn low-dimensional embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning [26, 53, 18], because of their superior performance across a broad spectrum of applications [58, 62, 63]. Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact [10, 47, 9, 50]. Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
Hyperdimensional computing [22, 23, 37, 49] offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector [22, 23]. All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products [22, 37, 49]. Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware [27, 28]. Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning [37, 28, 15, 34].
Hyperdimensional encodings of graphs have been investigated by Poduval et al. [43] and Nunes et al. [38]. These works focus on performing graph operations, e.g., similarity based retrieval, graph classification, etc. In contrast, our focus is on methods that learn node-level representations for node classification and link prediction.
Hyperdimensional node embeddings have been explored by Kang et al. [24] and [31]. Specifically, Kang et al. [24] proposed RelHD, a novel processing-in-memory (PIM) hardware architecture [35] based on FeFET technology [4]; and Li et al. [31] introduced HyperNode for learning HD embeddings of nodes. However, both RelHD and HyperNode suffer from significant shortcomings associated with how they construct node representations from node features (see section 3.5 for details).
1.2 Overview and Contributions
We introduce a novel method for transductive learning on graphs using hyperdimensional representations. The proposed approach encodes data samples using random projections into a very high- dimensional space (hyperdimensional or HD space for short) [22, 23, 37, 49]. It obviates the need for expensive iterative training of the sort required by deep learning methods. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
Contributions. The key contributions of the paper are as follows:
  • We introduce HDGL, a method for learning HD node embeddings that match the expressive power of GNN embeddings. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
  • We show how the HD node embeddings learned by HDGL can be used to perform both semi-supervised node labeling and link prediction tasks. Thus, unlike deep learning methods usually require separate models for each of these tasks, HDGL produces a single model that can support both tasks.
  • We present results of extensive experimental comparisons of HDGL with the state-of-the-art GNN methods on several standard benchmark data sets which show that HDGL achieves performance that is competitive with that of GNN, while avoiding the need for computationally expensive training.
  • We present results experiments which show that HDGL offers a better alternative to graph neural networks for settings that call for class-incremental learning, where the model is required to effectively discriminate between a growing number of classes that are presented incrementally over time.
Organization. The rest of the paper is organized as follows: Section 3 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN. Section 4 introduces HDGL, the proposed HD computing based algorithm for learning node embeddings. Section 5 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods. Section 6 concludes with a brief summary of the paper and an outline of some directions for further research.
2 Problem Statement
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$.
For the node label prediction problem in the transductive setting, let $y_{w}\in{1,\cdots,L}$ represent the label associated with each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$. This implies that not all nodes in the graph are labeled. The goal is to learn a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, the feature vector $\mathbf{x}_{v}$, and the labels $y_{w}$, to accurately predict the labels for a set of nodes that are unlabeled in the graph $\mathcal{G}$.In the transductive link prediction setting, the goal is to infer missing edges within the graph by leveraging the latent representations $\phi(v)$ of the nodes.
3 Background and Preliminaries
3.1 Properties of HD Computing
Hyper-dimensional computing, originally introduced by Kanerva [22] is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large [22, 23]. All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products [22, 37]. The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations [27, 28]. In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$. We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ [22]. This is also known as a blessing of HD spaces.
High Memory Capacity of HD Representation. As noted by Neubert et al. [37], in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits [1, 37, 49]. Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms [37].
3.2 HD Computing Operations
Bundling, denoted by $\oplus$, is a bitwise majority operation represented by :$\oplus:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. Bundling two randomly sampled HD vectors results in a new vector which is similar to the input vectors. Bundling $\oplus$ is associative and commutative i.e $\mathbf{a}\oplus(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\oplus\mathbf{b})\oplus\mathbf{c}=(\mathbf{c}\oplus\mathbf{a})\oplus\mathbf{b}$. Note that in case of Bundling an even number of HD vectors, there is a possibility of ties since Bundling is a bitwise majority operation. To break ties, a tie-breaking policy, e.g., random choice of 0 or 1 on dimensions where ties occur. Bundling can be thought of as the HD space analog of the mean operation euclidean spaces. Bundling, because it commutes, can be used to represent sets as well as multisets of objects encoded using HD vectors.
Binding, denoted by $\otimes$, is a bitwise Exclusive OR operation represented by :$\otimes:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, is an operation which takes two input vectors and maps them to a new vector. An important property of Binding is that Binding two HD vectors results in a vector that is dissimilar to both. Formally, let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two sampled HD vectors and $\mathbf{c}=\mathbf{a}\otimes\mathbf{b}$. Due to the near-orthogonality of HD vectors, we can then say that $d_{H}[\mathbf{c},\mathbf{a}]\approx 0.5$ and $d_{H}[\mathbf{c},\mathbf{b}]\approx 0.5$. Binding is invertible, associative and commutative i.e $\mathbf{b}\otimes(\mathbf{a}\otimes\mathbf{b})=(\mathbf{b}\otimes\mathbf{a})\otimes\mathbf{b}=\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$. Moreover, Binding is distributive over bundling i.e $\mathbf{a}\otimes(\mathbf{b}\oplus\mathbf{c})=(\mathbf{a}\otimes\mathbf{b})\oplus(\mathbf{a}\otimes\mathbf{c})$. Note that Binding also has a reflective property i.e the distance between two vectors doesn’t change when both vectors are bound to the same vector. Formally, $d_{H}[\mathbf{a},\mathbf{b}]=d_{H}[\mathbf{c}\otimes\mathbf{a},\mathbf{c}\otimes\mathbf{b}]$. Note that this also works even vectors are slightly different i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\mathbf{c}\otimes\mathbf{a}^{\prime},\mathbf{c}\otimes\mathbf{b}]$ if $\mathbf{a}\approx\mathbf{a}^{\prime}$.
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler [14] in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
Note that the operations referenced earlier have equivalents in the bipolar space $\{-1,1\}^{n}$. Specifically, the Bundle operation can be represented as signed addition in the bipolar space, while the binding operation aligns with multiplication in the bipolar space.
3.3 GNN-based Node Embdeddings
[…]
3.4 Representational Power of GNN
Several authors have investigated the representational power of GNN, including their ability to distinguish between different graph structures [59], e.g., using the Weisfeiler-Lehman (WL) graph isomorphism test [56, 19] , or to learn a set of functions on graphs [44], or the equivalence of the preceding two notions of expressivity[8]. The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors [56, 2]. The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors [19]. Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into distinct encodings. As shown by [59, 33], GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations [59].
3.5 Limitations of Previous HD Graph Methods
The HD Computing based graph-level learning methods introduced by Nunes et al. [38] and Poduval et al. [43] are designed for similarity-based graph retrieval and graph classification, as opposed to node classification or link prediction. Nunes et al. [38] constructs the Graph HD vector using node features derived using PageRank [5]. Poduval et al. [43] assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector. Thus, both methods ignore node features present in the graph.
RelHD [24] constructs node representations for node classification tasks and uses the resulting representation to design a novel hardware chip design with a scalable processing in-memory (PIM) architecture based FeFET. However, RelHD has a significant limitation with respect how it encodes the node features into HD vectors. Specifically, it treats node features as dimension-wise indicator variables and computes the HD vector representation of each node by first generating $d$ random basis vector. Let $\Upsilon_{p}$ be such a random basis Hyper vector. Random basis vectors $\Upsilon_{p}$ are bundled/accumulated if feature at index $p$ is present i.e HD vector for a node $u$ is computed using basis vectors and raw features $\mathbf{x}_{u}=[x_{u1},\ldots,x_{ud}]$ using the following formula:- $\oplus_{p\in\{1,\ldots,d\}}(\mathbbm{1}(x_{up}\neq 0)\cdot\Upsilon_{p})$. Consequently, this method is not suitable for graphs with continuous or integer-valued node attributes in continuous form or even simple bag-of-words counts. Note that because HyperNode [31] also utilizes the same mechanism for constructing HD vector encodings of the nodes, it suffers from the same limitations as RelHD.
Furthermore, none of the aforementioned methods [38, 43, 24, 31] support link prediction on graphs.
In contrast, HDGL offers an effective and efficient method for constructing more expressive node embeddings that can accommodate not only binary or multi-valued, but also integer-valued or continuous node features. The resulting HD encodings of nodes support both node classification and link prediction.
4 Hyper-Dimensional Graph Learner
In essence, our design of HDGL aims to achieve a high level of expressive power similar to graph neural networks (GNN). We utilize HD computing operations on node representations, taking advantage of HD operations—Bundling ($\oplus$), Binding ($\otimes$), and Rotation ($\Pi$) —and their properties. This ensures that HDGL, akin to GNN, not only aspires to build expressive power but also learns to map similar graphs to similar representations.
The proposed HD learner carries out 3 major steps: (i). Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$. (ii). Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$. (iii). Use $\mathbf{z}_{v}$ to predict $y_{v}$ or predict links/edges in the graph. We proceed to describe each of these steps in detail.
Figure 1: An example mapping node features to a binary space using Random Hyperplanes. A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side. Utilizing multiple hyperplanes results in a binary sketch for each point.
4.1 Mapping Node Features to HD-space
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations [12], a type of Locality Sensitive Hashing [21, 7]. The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$. The primary purpose of adding the $\Gamma$ offset is to ensure that the not all sampled hyperplanes pass through the origin. When dealing with evenly spread points, typically in the context of real-valued vectors, setting $\Gamma$ to zero is often appropriate.
Note that the resulting binary sketch of each node is in fact a $\beta$-dimensional binary HD vector. Nodes with similar features will have similar HD representations; and nodes with dissimilar features will have dissimilar HD representations.
4.2 Neighborhood Aggregation Scheme
Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
Bundling for generalizing across similar nodes. We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets). Hence, we propose to use Bundling to realize $\varphi_{k}$. More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective. However, as noted by [59], GNNs that use the mean to aggregate features over node neighborhoods [26, 18] perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
It is worth noting that the Binding operation $\otimes$ would be inappropriate in this context because binding of two similar vectors results in dissimilar vectors. Thus, if Binding were used to realize $\varphi_{k}$, similar yet distinct local neighborhoods structures would yield different HD representations, adversely impacting HDGL’s ability to generalize across such graph structures. To complicate matters, Binding is also invertible i.e $\mathbf{a}\otimes\mathbf{b}\otimes\mathbf{b}=\mathbf{a}$; and hence, it could discard information that is useful for downstream task e.g., node label prediction.
Binding for linking node with its neighborhood. Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
We propose to use Binding the operation $\otimes$ to realize the function $\phi(\cdot)$ because it offers a means to balance the discriminative power of the resulting model against its generalization ability. We illustrate the preceding claim with an example. Let $v\in\mathcal{V}$ and $w\in\mathcal{V}$ be two nodes in a graph. Suppose $\phi(\cdot)$ is realized using the Binding Operation $\otimes$ and $\psi_{k}(\cdot)$ is realized using the Bundling Operation. Considering only 1-hop and 2-hop neighbors of $v$, the latent representation of $v$ and $w$ will be
$\begin{split}\mathbf{z}_{v}&=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}&=\mathbf{r}_{w}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(w)}\mathbf{r}_{m}\Big{)}\end{split}$ (1)
Suppose we have
$\begin{split}\mathbf{r}_{w}\approx\mathbf{r}_{v};\quad\;\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\approx\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(w)}\mathbf{r}_{j}\Big{)}\\ \end{split}$
and ${\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}}$ is dissimilar to ${\mathop{\oplus}_{j\in\mathcal{N}^{2}(w)}\mathbf{r}_{j}}$. Then $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since binding is reflecting i.e
$d_{H}\Big{[}\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)},\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(w)}\mathbf{r}_{j}\Big{)}\Big{]}\approx d_{H}\Big{[}\mathbf{z}_{v},\mathbf{z}_{w}\Big{]}$
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired. On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
Conversely, it is easy to see that Bundling is not a good choice for realizing $\phi(\cdot)$. Bundling of the node vector $\mathbf{r}_{v}$ with neighborhood information vector $\oplus_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$, a vector that has already been bundled with $k$-hop neighbors, will result in a vector that predominantly represents information from a node’s neighborhood, while leaving out essential information about the itself. This is due to the fact that Bundling is a bitwise majority operation; and hence, if bundling is used to realize $\phi(\cdot)$, the information from a node’s neighborhood suppresses the information about the node itself.
Table 1: Dataset Summary.
Dataset Nodes Edges Features Feature Type Classes Label Rate Edge Density
CORA 2708 5429 1433 $\{0,1\}$ 7 0.052 0.0004
CiteSeer 3327 4732 3703 $\{0,1\}$ 6 0.036 0.0004
PubMed 19717 44338 500 $(0,1)$ 3 0.003 0.0001
BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$ 6 0.023 0.0127
Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$ 15 0.016 0.0001
Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$ 5 0.002 0.0001
DBLP 17725 105781 768 $\mathbb{R}$ 4 0.004 0.0006
Improving injectivity via Rotation. Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as $\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}$. Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm. To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information. To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
where $\Pi(\cdot)$ and $\Pi\;\Pi(\cdot)$ denote single and double rotation operations respectively. The preceding equation ensures that $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ will be dissimilar since rotation of a HD vector yields a HD vector that is dissimilar/orthogonal to the original. That is,
$\begin{split}d_{H}\Big{[}\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)},\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\Big{]}\approx 0.5\\ d_{H}\Big{[}\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\big{)},\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\Big{]}\approx 0.5\\ \end{split}$
Hence, considering only 1-hop and 2-hop neighbors, the proposed aggregation scheme to be used by HDGL for computing the HD representation of a node $v\in\mathcal{V}$ in the graph is given by
$\begin{split}\mathbf{z}_{v}=\mathbf{r}_{v}\otimes\Pi\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{1}(v)}\mathbf{r}_{j}\Big{)}\otimes\Pi\;\Pi\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{2}(v)}\mathbf{r}_{m}\Big{)}\\ \end{split}$ (2)
4.3 Node Label Prediction using HDGL
With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects [15]. Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows: $\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}]$.
4.4 Link Prediction Using HDGL
Section 4.1 and 4.2 outline a technique for constructing node representations optimized for node classification. Here we show that the same node representation can be used to support the transductive link prediction task.
Kanerva [22, 23] describes a method for encoding entries of a dataset. Let $\mathbf{a}_{1}$, $\mathbf{a}_{2}$, $\mathbf{a}_{3}$ be some randomly sampled HD vectors representing data entry attributes (eg. weight, height and gender). Let $\mathbf{\alpha}$, $\mathbf{\beta}$ and $\mathbf{\gamma}$ be values for the corresponding attributes represented in HD format. Let $\mathbf{r}$ be a HD vector such that: $\mathbf{r}=(\mathbf{a}_{1}\otimes\mathbf{\alpha})\oplus(\mathbf{a}_{2}\otimes\mathbf{\beta})\oplus(\mathbf{a}_{3}\otimes\mathbf{\gamma}).$ From the properties of HD-computing [22], it is possible to retrieve values for a corresponding attribute using the Binding $\otimes$ operator. Specifically, $\mathbf{\alpha}\approx\mathbf{r}\otimes\mathbf{a}_{1};\quad\mathbf{\beta}\approx\mathbf{r}\otimes\mathbf{a}_{2};\quad\mathbf{\gamma}\approx\mathbf{r}\otimes\mathbf{a}_{3};$
We apply the same concept to link prediction. Let $\mathbf{e}^{+}$ be a hypervector storing information about all or some of the edges present in the graph, and let $\mathbf{e}^{-}$ be a hypervector representing information about some of the non-edges in the graph. Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Let $\mathbf{Z}$ have $N$ rows, storing latent representations of all the nodes in the graph. That is, a row in $\mathbf{Z}$ is represented by $\mathbf{z}_{u}$, where $u\in\mathcal{V}$. Now that we have encoded edges and non-edges in the graph, the goal is to find new or missing links in the graph. This can also be thought of as reconstructing/generating the adjacency matrix.
To find possible new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph. This results in a new matrix $\mathbf{Z}^{+}$ where each row is $\mathbf{z}_{u}\otimes\mathbf{e}^{+}$ where $u\in\mathcal{V}$. $\mathbf{Z}^{+}$ represents the possible node representations of the target nodes in a link, with the corresponding source node representation being $\mathbf{Z}$. We follow the same procedure for the non-edge hypervector, resulting in $\mathbf{Z}^{-}$. Precisely:
$\mathbf{Z}^{+}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}$ $\mathbf{Z}^{-}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}$
Now, we find pairwise normalized distances between rows of $\mathbf{Z}^{+}$ and rows of $\mathbf{Z}$. This results in $\mathbf{D}^{+}=dist(\mathbf{Z}^{+},\mathbf{Z})\in[0,1]^{N\times N}$ , a matrix representing distances. If entry $(i,j)$ is close to zero, it indicates that the distance between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting that $(i,j)$ might be a possible link. We perform the same process for $\mathbf{Z}^{-}$, resulting in the $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})\in[0,1]^{N\times N}$ matrix. However, an entry close to zero in $\mathbf{D}^{-}$ indicates a high possibility of not having an edge between the corresponding nodes.
We now construct a prediction of the adjacency matrix represented by $\hat{\mathbf{A}}$, where entries $\hat{A}_{ij}\in[0,1]$. To construct the final reconstructed or predicted adjacency matrix of probabilities (based on distances), which factors in both the distances of having, or not having an edge using $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$, we compute:
$\hat{A}_{ij}=\begin{cases}\sigma\big{(}(1-D^{+}_{ij})+D^{-}_{ij}\big{)}&\text{if }D^{+}_{ij}<D^{-}_{ij}\\ \sigma\big{(}D^{+}_{ij}-(1-D^{-}_{ij})\big{)}&\text{else}\end{cases}$
where $\sigma(\cdot)$ is the sigmoid function. The rationale behind this operation lies in the comparison between $D^{+}_{ij}$ and $D^{-}_{ij}$. When $D^{+}_{ij}<D^{-}_{ij}$, it indicates a higher likelihood of a connection between nodes $i$ and $j$ according to the distance metric. Hence, we apply the operation $\sigma((1-D^{+}_{ij})+D^{-}_{ij})$ to obtain a value closer to one. We use the same logic when $D^{+}_{ij}\geq D^{-}_{ij}$ to obtain a value close to zero.
5 Experiments
We proceed to report results of extensive experiments comparing HDGL with a set of state-of-the-art node classification methods and link prediction methods, including several GNN models, in terms of both predictive performance and runtime.
5.1 Data sets
Since HDGL is designed in such a way that nodes with similar features and neighborhoods will have similar representations, we exclusively concentrate on homophilous graphs in this study, deferring the exploration of more complex cases involving heterophily graphs to future research.
For our experiments, we utilize commonly used node labeling homophilous graph benchmarks: the CORA graph dataset [61], sourced from McCallum et al. [32]; the CiteSeer graph dataset [61], sourced from Giles et al. [16]; and the PubMed dataset [36]. Additionally, we include three more datasets from the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. [45], and the BlogCatalog Graph dataset from Huang et al. [20]. Additionally, we introduce a new version of the DBLP dataset, built upon the works of Pan et al. [39]. We remove nodes with no neighbors in the graph obtained from Pan et al. [39]. The original node features from Pan et al. [39] consist of manuscript titles. We utilize BERT [11] to obtain embeddings for the titles and employ them as node features. The data sets and their statistics are given in Table 1.
We adhere to the Train/Validation/Test splits described in Yang et al. [61], Kipf and Welling [26], and Veličković et al. [53] for the CORA, CiteSeer, and PubMed datasets, respectively. For the Coauthor CS, Coauthor Physics, BlogCatalog, and DBLP datasets, we employ the data splitting strategy outlined in Shchur et al. [45].
We test link prediction models on CORA, CiteSeer, and Pubmed datasets. These models are trained on incomplete data, retaining node features but omitting some edges. Validation and test sets include excluded edges and randomly selected unconnected node pairs. Evaluation assesses the models’ ability to classify both edges and non-edges. We follow splitting strategy from Kipf and Welling [25] work i.e Validation and test sets each contain 5% and 10% of citation links.
5.2 Models
For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk [42] and Label Propagation [64]. We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) [26], Graph Attention Network (GAT) [53], and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) [57] which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input. We also assess our model against RelHD [24], an existing HD algorithm used for node classification.
For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk [42] and Spectral clustering (SC) [48]. We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling [25].
5.3 HDGL Model Specifications
To ensure that HDGL are directly comparable with most GNN architectures which typically use 2 layers, and hence aggregate information over 2-hop neighbors, we limit HDGL to aggregate information over at most 2-hops i.e HD node representations are calculated as per Eq.(2).
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog. For BlogCatalog we sample 21 of the 1-hop neighbors and 51 of the 2-hop neighbors since Blogcatalog graph is denser than the other datasets.
For node classification in CORA, CiteSeer, and BlogCatalog datasets, we map node features to $50,000$-dimensional HD vectors through random projections. Similarly, for the PubMed, Coauthor CS, Coauthor Physics, and DBLP datasets, we map node features to $20,000$-dimensional HD vectors using the same method. While these choices may seem arbitrary, it’s crucial to note that selecting dimensions between $20,000$ and $50,000$ typically suffices to maintain the desired properties of HD-computing. This assertion finds extensive support in the research from Kanerva [22, 23].
Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features. In these datasets, RelHD performs comparably to HDGL)
Method Cora Citeseer Pubmed BlogCat CompSci Physics DBLP
LogReg 52.2 $\pm$ 0.5 46.3 $\pm$ 0.4 67.9 $\pm$ 0.5 65.3 $\pm$ 1.0 86.8 $\pm$ 0.6 88.2 $\pm$ 0.9 56.2 $\pm$ 3.0
DeepWalk 65.3 $\pm$ 1.2 46.2 $\pm$ 0.9 69.3 $\pm$ 0.9 58.6 $\pm$ 1.3 78.1 $\pm$ 0.9 86.4 $\pm$ 0.9 66.1 $\pm$ 2.2
Label Prop 70.9 $\pm$ 0.0 47.7 $\pm$ 0.0 71.3 $\pm$ 0.0 47.8 $\pm$ 3.8 76.6 $\pm$ 0.9 85.6 $\pm$ 1.1 66.4 $\pm$ 1.6
GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 69.5 $\pm$ 0.4 91.0 $\pm$ 0.7 92.3 $\pm$ 0.6 72.1 $\pm$ 0.8
GAT 82.8 $\pm$ 0.5 72.0 $\pm$ 0.6 79.1 $\pm$ 0.3 65.3 $\pm$ 1.2 90.9 $\pm$ 0.6 92.5 $\pm$ 0.9 75.4 $\pm$ 1.0
SGC 79.8 $\pm$ 0.5 68.5 $\pm$ 0.4 75.4 $\pm$ 0.5 70.8 $\pm$ 0.3 89.1 $\pm$ 0.6 92.1 $\pm$ 0.7 76.3 $\pm$ 0.9
RelHD 78.9 $\pm$ 0.5 69.3 $\pm$ 0.4 69.1 $\pm$ 0.4 60.9 $\pm$ 1.2 85.5 $\pm$ 0.8 85.3 $\pm$ 0.9 44.9 $\pm$ 0.2
HDGL 79.5 $\pm$ 0.7 70.0 $\pm$ 0.6 76.8 $\pm$ 1.1 69.3 $\pm$ 0.5 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9 68.7 $\pm$ 1.2
5.4 Training and Evaluation
Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters. Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance. In contrast, learning methods such as HDGL which invole no iterative training do not need validation data. Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question. Thus, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. For tuning the hyperparameters, we used grid search over a set of 12 configurations. In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay. In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored. Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4]. In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6]. GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3. The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three GNN models.
As LogReg, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propagation determines the balance between existing and propagated labels. Higher alpha prioritizes existing labels; lower alpha favors propagated ones. to 0.5. For DeepWalk, we utilize the DGL library, setting the embedding dimension to 32 and the number of walks to 20, with a batch size of 64. We sample such walks for all nodes in the graph and train DeepWalk model for 10 epochs and then employ logistic regression for classification, incorporating early stopping based on the validation set.
Following Kang et al. [24], the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
We report measured performance on the test data, averaged across 10 random selections of pertinent parameters. These parameters include the random hyperplanes employed for constructing high-dimensional representations used by HDGL, random basis vector initializations for RelHD, weight initializations of the deep learning models, and data splits for datasets such as Coauthor Physics, Coauthor CS, BlogCatalog, and DBLP. For Cora, Citeseer, and Pubmed, we use predefined splits given in Yang et al. [61].
For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling [25].
5.5 Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework [40], complemented by the Deep Graph Library (DGL) [54] for importing datasets and implementing deep learning models. None of our experiments used parallel execution or GPUs. We have also developed a PyTorch version of RelHD, as the codebase was not publicly available and experiments detailed in Kang et al. [24] were reliant on hardware architecture implementations. We undertake this effort to compare with our model and use RelHD as a baseline in this study.
5.6 Results
Node Classification Performance. The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp). The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms. This is particularly notable with datasets lacking $\{0,1\}$ features. However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
To highlight the advantage of HDGL more prominently, we conducted experiments for node classification on the CORA and CiteSeer datasets where we reduced the node features to 100 dimensions using PCA [41], thereby trying to enforce continuous node features. As depicted in Figure 2, the impact is evident through a noticeable drop in performance for RelHD. Conversely, for other methods including HDGL, the performance remains relatively unchanged.
Figure 2: Impact of performing PCA with a dimensionality of 100 on features for node classification accuracy.
Figure 3: Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations. Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
Table 3: Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling [25] paper.
Methods Cora Citeseer PubMed
AUC AP AUC AP AUC AP
DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
SC 0.846 0.885 0.791 0.826 0.849 0.888
GAE 0.910 0.920 0.895 0.899 0.964 0.965
VGAE 0.914 0.926 0.908 0.920 0.944 0.947
HDGL 0.849 0.880 0.768 0.842 0.853 0.884
Node Classification Run-times. Figure 3 compares the run-time for the node classification task between HDGL and RelHD (both of which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyperparameter tuning), alongside the run-time of HDGL. We observe that, the run-time of HDGL is substantially less than that of all the state-of-the-art GNNs. While in most cases, RelHD exhibits slightly or significantly worse run times compared to HDGL, it’s crucial to note that HDGL consistently delivers superior performance across various datasets when compared to RelHD.
Table 4: Class Incremental Node Labeling Results. No hyperparameter grid search is used for GCN and SGC algorithms. Wall-time includes both training and inference of the model. HDGL does not require retraining from $t=2$.
Model BlogCatalog Coauthor Physics
t = 1 t = 2 t = 3 t = 4 t = 5 t=1 t=2 t=3 t=4
GCN Accuracy 94.94 85.81 75.16 74.82 69.61 98.95 97.33 95.57 92.58
Runtime 51.2s 18.7s 16.5s 33.9s 23.9s 65.9s 62.3s 65.8s 51.2s
SGC Accuracy 94.56 82.31 73.23 72.72 69.81 98.94 97.39 95.36 91.81
Runtime 6.3s 3.72s 3.43s 3.73s 3.6s 9.1s 8.5s 8.8s 9.7s
HDGL Accuracy 89.11 88.49 75.56 71.73 69.91 98.56 94.87 92.53 90.8
Runtime 20.2s 0.3s 0.4s 0.4s 0.6s 38.3s 1.5s 1.7s 1.7s
It is crucial to note that while we have used grid search on 12 potential setups for the deep learning model hyperparameter optimization of deep learning models typically explores a much larger space. This expanded search naturally comes with increased runtime, especially when GPUs are utilized, further widening the energy efficiency gap between Graph Neural Networks (GNN) and traditional deep learning models like HDGL.
Link Prediction Performance. Even though HDGL representations are primarily tailored for transductive node classification, their performance for link prediction is comparable to that of DeepWalk and Spectral clustering; as seen from Table 3. However, they do not reach the same level of effectiveness as GAE and VGAE, which leverage GNN methods. Despite lacking iterative training like deep learning methods, HDGL still manages to achieve respectable results in link prediction. Moreover, unlike deep learning approaches that typically require distinct models for link and node classification, HDGL tackles both tasks within a unified framework. We also perform link prediction experiments by varying the dimensionality of the HDGL model. This exploration arises from observing a slight yet discernible variation in performance for link prediction tasks, unlike the more consistent outcomes observed in node classification experiments. As mentioned earlier, typically, dimensions ranging from 20,000 to 50,000 are sufficient for HD properties to hold. However, higher dimensionality generally leads to better performance. Nevertheless, beyond a certain threshold, the performance tends to plateau, as illustrated in Figure 4.
Figure 4: Link Prediction Performance of HDGL under Various dimensionality configurations.
HDGL in the Class-Incremental Learning Setting Consider a scenario with a dynamic graph where nodes and edges remain static, but the new node labels appear over time i.e initially, only a subset of nodes is labeled, belonging to two classes. Over time, new nodes receive labels; thus, introducing new classes. This is a node labeling instance of the class-incremental learning problem [52, 3, 51]. In this setting, deep learning methods typically require either retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes [51]. This retraining incurs significant additional runtime and computation.
In contrast to deep neural networks, which necessitate retraining, Hyperdimensional Graph Learner (HDGL) operates without an iterative training phase. Instead, it constructs label hypervectors using predefined latent node representations. This approach eliminates the need for retraining, resulting in substantial savings in both runtime and computation costs.
To simulate incremental node classification on a fully labeled graph, we use BlogCatalog and Coauthor Physics Dataset. We first split the nodes into train/val/test as in Shchur et al. [45]. We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are only available; and so on such that at the final time step, all of the labels $\{1,\cdots,L\}$ are available.
To simplify matters for GNN models, we assume that the models are provided the list of $L$ possible labels although only a subset have been encountered in the training data. This allows the last layer of the GNN to have $L$ outputs. Also, we forego hyperpameter search for the GNN models to minimize the computational overhead.
In comparing HDGL against GCN and SGC in our experimental setup, we opted not to include the GAT model due to its longer runtime, as depicted in Figure 3. As seen from Table 4, we see that HDGL incurs a significant computational overhead only at timestep 1, necessitating the computation of node representations followed by inference. However, for subsequent timesteps , HDGL leverages previously computed node representations, requiring only inference. In contrast, deep learning models like GCN and SGC mandate retraining for each timestep to accommodate new labels, resulting in comparatively slower runtimes.
6 Conclusion
We introduced HDGL, a novel and efficient algorithm for transductive learning on graphs using hyperdimensional representations. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks. HDGL obviates the need for expensive iterative training of the sort required by deep learning methods. HDGL is also well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
Our experiments with widely used benchmark datasets show that on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost. Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods. We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.